Suppose that during the past year, only one of 100 similar plants owned by Hauser Co. suffered an internal fire. With 125 employees in the plant at the time, 14 casualties resulted from the fire and 2 of these 14 were fatalities (died). Based on this information, answer the following questions:

A      What is the probability of a fire at any given plant during the upcoming year? (8 points)

B      What is the probability of a casualty (injured or died), given that there is a fire? (8 points)

C      What is the probability of a fatality that results from a fire in a plant during the upcoming year? (joint probability) (9 points)
Answer #1

a) probability of a fire at any given plant during the upcoming year =1/100 =0.01 (as out of 100; only one had fire)

b) probability of a casualty (injured or died), given that there is a fire =14/125=0.112 (as 14 out of 125 got casuality)

c)

probability of a fatality that results from a fire in a plant during the upcoming year

=2/125 =0.016